OpenStudy (anonymous):

what is the solution set; use the multiplication principle. -6x>1/13

OpenStudy (saifoo.khan):

\[x < \frac{1}{-6*13} \]

OpenStudy (saifoo.khan):

\[x < \frac{1}{-78} \]

OpenStudy (anonymous):

so the top number always stays the same?

OpenStudy (saifoo.khan):

\[(\frac{1}{-78},\infty)\]

OpenStudy (saifoo.khan):

No, that isn't important, when we cross multiply, this happens.

OpenStudy (anonymous):

what would the inequality symbol be?

OpenStudy (saifoo.khan):

you want the solution set right, so that is going be this way!

OpenStudy (anonymous):

this is what {xIx__ __}

OpenStudy (anonymous):

*it gives me

OpenStudy (saifoo.khan):

If its this way then, \[x|x<\frac{-1}{78}\]

OpenStudy (anonymous):

oh ok. I did the right way then. Thank you

OpenStudy (saifoo.khan):

Cool.